40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13032     Entry time: Fri Jun 2 00:54:08 2017     Reply to this: 13033
Author: Koji 
Type: Update 
Category: ASS 
Subject: Xarm ASS restoration work 

While Gautam is working the restoration of Yarm ASS, I worked on Xarm.

Basically, I have changed the oscillator freqs and amps so as to have linear signals to the misalignment of the mirrors.
Also reduced the complexity of the input/output matrices to avoid any confusion.

Now the ITM dither takes care of the ITM alignment, and the ETM dither takes care of the ETM alignment.
The cavity alignment servos (4dofs) are running fine although the control band widths are still low (<0.1Hz).
The ETM spot positions should be controlled by the BS alignment, but it seems that these loops have suspicion about the signal quality.

While Gautam wa stouching the input TTs, we occasionally saw anomalously high transmission of the arm cavities (~1.2).
We decided to use this beam as this could have indicated partial clipping of the beam somewhere in the input optics chain.

Then the arm cavity was aligned to have reasonably high transmission for the green beam. i.e. Use the green power mon PD as a part of the alignment reference.

This resulted very stable transmission of both the IR and green beams. We liked them. We decide to use this a reference beam at least for now.

Attachment1: GTRX image at the end of the work.

Attachment2: ASSX screen shot

Attachment3: ASSX servo screen shot

Attachment4: Green ASX servo screen shot

Attachment 5: Screen shot of the ASS X strip tool

Attachment 6: Screen shot of the ASS X input matrix

Attachment 7: Screen shot of the ASS X output matrix

Attachment 1: GTRX.jpeg  18 kB  Uploaded Fri Jun 2 01:55:30 2017  | Hide | Hide all
GTRX.jpeg
Attachment 2: 54.png  70 kB  Uploaded Fri Jun 2 02:09:41 2017  | Hide | Hide all
54.png
Attachment 3: 37.png  22 kB  Uploaded Fri Jun 2 02:10:31 2017  | Hide | Hide all
37.png
Attachment 4: 16.png  16 kB  Uploaded Fri Jun 2 02:10:34 2017  | Hide | Hide all
16.png
Attachment 5: 26.png  21 kB  Uploaded Fri Jun 2 02:11:03 2017  | Hide | Hide all
26.png
Attachment 6: 41.png  5 kB  Uploaded Fri Jun 2 02:12:05 2017  | Hide | Hide all
41.png
Attachment 7: 01.png  4 kB  Uploaded Fri Jun 2 02:12:07 2017  | Hide | Hide all
01.png
ELOG V3.1.3-